### v1

# Nullifiers of non-Gaussian cluster states through homodyne measurement

Abstract
Summary: Multiplexing potential of large-scale optical Gaussian cluster states was already experimentally demonstrated. Non-Gaussian states must be embedded in a cluster state to reach universality and fault tolerance. Basic quantum operations with the potential to induce non-Gaussianity, such as photon subtraction, start to broaden the available experimental toolbox as well. The interplay of the two, non-Gaussian multimode cluster states, remains challenging both theoretically and experimentally. Here we propose a framework for characterization of non-Gaussian cluster states. In analogy to nowadays frequently used Gaussian nullifiers, we define non-Gaussian nullifiers of cluster states that enable certification of non-Gaussianity in the cluster state. We illustrate the concept on an example of cluster state created from photon subtracted squeezed states. The nullifier can be evaluated from homodyne measurement, which is a natural detection for optical cluster states.
The main advantage of quantum optics as a platform for continuous variable quantum computing Menicucci (2014); Larsen et al. (2021); Aghaee Rad et al. (2025) is the scalability of the number of bosonic modes. Deterministic entanglement generation via Gaussian operations enables feasible preparation of enormously large entangled states called cluster states van Loock et al. (2007). Those have already been experimentally demonstrated multiple times Larsen et al. (2019); Asavanant et al. (2019); Yoshikawa et al. (2016). The great remaining challenge to unlock the universality and fault tolerance on this platform is the preparation of non-Gaussian states with sufficient quality and complexity and their involvement in the cluster state Gu et al. (2009); Konno et al. (2024); Kala et al. (2024).
The experimental preparation of continuous variable Gaussian cluster states consists of a set of squeezed states that are entangled via beam splitter operations Larsen et al. (2019); Asavanant et al. (2019). Individual modes of the generated cluster states are eventually measured via homodyne detection that enables its characterization. Squeezing of noise in mutually commuting operators called nullifiers shows that the experimental implementation approaches approximation of the ideal cluster state van Loock et al. (2007). An ideal cluster state is a ground state of the nullifiers, leading to their simultaneously vanishing noise. The experimental version, burdened by decoherence and finite squeezing, is then certified when its noise in the nullifiers is lower than a given threshold Larsen et al. (2019).
On the other hand, experimental preparation of non-Gaussian states is far more challenging as the non-Gaussian features can be more fragile than Gaussian squeezing in the presence of decoherence and imperfections Jeannic et al. (2018); Kala et al. (2022). The most common approach to introduce non-Gaussianity into light is via photon-resolving measurement in the Fock basis, performed on a part of an entangled state. This principle lies at the core of many state preparation protocols Konno et al. (2021, 2024); Fadrný et al. (2024). Certifying successful preparation of non-Gaussian state is a complex task. Partially due to their fragility, but also due to the many forms they can attain Ogawa et al. (2016); Ourjoumtsev et al. (2006); Asavanant et al. (2017); Fadrný et al. (2024); Konno et al. (2024). Properties shared by all non-Gaussian states, like Wigner negativity, often lack a straightforward operational interpretation in quantum information applications. Quality of experimental realization of a specific target state can be evaluated via excess noise carried in some operator that has the ideal version as a ground state Konno et al. (2021); Kala et al. (2022); Konno et al. (2024); Marek (2024); Kala et al. (2025); Šimon Bräuer et al. (2025). The non-Gaussianity is then certified by comparison to a minimal noise attainable within Gaussian states Miyata et al. (2016); Konno et al. (2021); Kala et al. (2022); Konno et al. (2024).
The successful implementation of continuous variable quantum computing then lies in embedding the non-Gaussian states into the cluster state. Primarily, involving non-Gaussian states is necessary to avoid efficient classical simulation enabled by Gaussian systems Chabaud and Walschaers (2023). To this end, photon subtraction or addition can be used to prepare multimode non-Gaussian states Walschaers et al. (2017); Kala et al. (2025). Further, appropriate auxiliary states, like the cubic phase state, can be entangled with or prepared within the cluster state to unlock universality via measurement-induced nonlinear operations Miyata et al. (2016); Gu et al. (2009); Kala et al. (2024). Eventually, using GKP encoding of qubits into continuous-variable modes Gottesman et al. (2001) enables correcting for errors caused by finite squeezing in the cluster state Larsen et al. (2021). All mentioned approaches require entangling the cluster state with non-Gaussian states, i.e., preparing a non-Gaussian cluster state. Here, we present nullifiers for non-Gaussian cluster states. The concept enables certification of non-Gaussianity presence in the cluster state and can be evaluated from homodyne measurement.
I Non-Gaussian nullifiers
The canonical form of the Gaussian cluster state is prepared from eigenstates of the $p$ quadrature via continuous-variable $Cz$ gates Zhang and Braunstein (2006). This two mode gate transforms the $p$ quadratures as
$\begin{split}p_{1}\rightarrow&p_{1}+x_{2}\\ p_{2}\rightarrow&p_{2}+x_{1},\end{split}$ (1)
leaving the $x$ quadratures intact. The ideal Gaussian cluster state is found to be the ground state of the nullifiers
$N_{i}=p_{i}-\sum_{j\in N(i)}x_{j},$ (2)
which are a linear combination of quadrature operators. The summation goes over neighboring connected modes. The form is an inverse operation to the $Cz$ gates in the cluster state preparation and addresses the infinite squeezing of the $p$ quadrature eigenstates. An approximation created from finitely squeezed states then carries some noise in the nullifier operators.
In quantum optics, passive linear beam splitters are used instead of the $Cz$ gates as the latter require an active operation of online squeezing in their optical implementation van Loock et al. (2007). The preparation of the cluster state can be described in the Heisenberg picture as
$^{\prime}=V\textbf{r},$ r , (3)
where $\textbf{r}=(x_{1},p_{1},x_{2},p_{2}...)^{T}$ is a set of quadratures corresponding to vacuum states, the symplectic transformation $V$ describes the action of squeezers and $Cz$ gates, and $\textbf{r}^{\prime}$ quadratures presented for the homodyne measurement. The transformation $V$ can be decomposed via Bloch-Messiah decomposition into a sequence of passive linear optical network, a set of squeezer, and another passive linear network Braunstein (2005); van Loock et al. (2007). As the input consists of a vacuum state, the first linear optical network is irrelevant, and the whole cluster state preparation can be composed of a set of squeezers followed by a passive optical network with symplectic transformation $M$ van Loock et al. (2007). In this manner, the structure of the cluster state is created via a passive linear interferometer. Spatial-temporal multiplexing enables making the large cluster preparation compact Menicucci (2011).
Let us describe a two mode cluster state shown in Fig. 1 as an illustration. The cluster state has two nodes and a connecting edge. According to (2), the nullifiers are of the form
$\begin{split}N_{1}=&p_{1}^{\prime}-x_{2}^{\prime}\\ N_{2}=&p_{2}^{\prime}-x_{1}^{\prime}.\end{split}$ (4)
The preparation of the clusters state starts with squeezed states with quadrature operators $x_{1},p_{1},x_{2}$ and $p_{2}$ that are connected to vacuum via $x_{i}=\exp(-r)x_{i}^{(0)}$ and $p_{i}=\exp(-r)p_{i}^{(0)}$ for $i=1,2$. The passive linear optical network is shown in Fig. 1b, consists of phase shift applied to the second mode, beam splitter interaction and second phase shift on the second mode. The overall transformation yields
$\begin{split}x_{1}^{\prime}=&\frac{1}{\sqrt{2}}(x_{1}+p_{2}),\;p_{1}^{\prime}=\frac{1}{\sqrt{2}}(p_{1}-x_{2})\\ x_{2}=&\frac{1}{\sqrt{2}}(x_{2}+p_{1}),\;p_{2}=\frac{1}{\sqrt{2}}(p_{2}-x_{1}).\end{split}$ (5)
Which gives the nullifiers as
$\begin{split}N_{1}=&\sqrt{2}p_{1}=\sqrt{2}p_{1}^{(0)}\exp(-r)\\ N_{2}=&\sqrt{2}p_{2}=\sqrt{2}p_{2}^{(0)}\exp(-r),\end{split}$ (6)
thus vanishing as the initial squeezing goes to infinity $r\rightarrow\infty$
To reach universality and fault tolerance, it is necessary to embed non-Gaussian states into the Gaussian cluster state. In this case, the nullifiers of Gaussian cluster states can no longer be used for the cluster state certification in experimental data. We consider the following model of a cluster state. Some of the input modes are populated by a single-mode non-Gaussian states |$\ket{\psi}_{n}$, the rest are initialized in the vacuum state. The cluster state is prepared from the set of initial states via Gaussian squeezers followed by a linear interferometer with unitary operation $U$ as
|$\ket{\Psi}=U(\otimes_{n}\ket{\psi}_{n}\otimes_{k}\ket{0}).$ (7)
Further, we define the characteristics of the individual states |$\ket{\psi}_{n}$. Let $O_{n}$ be a positive semidefinit operator that has the state |$\ket{\psi}_{n}$ as a ground state, i.e. eigenstate of the lowest - zero eigenvalue, and let it have a representation by quadrature operators $x_{n}$ and $p_{n}$
$O_{n}=O_{n}(x_{n},p_{n}).$ (8)
Then the $k_{th}$ non-Gaussian nullifier can be defined as
$O_{k}=O_{k}(\{M^{-1}\textbf{r}^{\prime}\}_{k},\{M^{-1}\textbf{r}^{\prime}\}_{k+1}),$ (9)
where {}$\{\}_{k}$ denotes the $k$-th row and $M$ is a symplectic transformation equivalent to the evolution generated by the unitary operator U. The inverse symplectic transformation enables us to untwist the cluster state structure and address the statistics of the initial modes. Moreover, as the cluster was created by a linear operation, the ”untwisting” can be done on the measured data - a technique already used in the current verification of Gaussian cluster state preparation.
An experimentally prepared non-Gaussian cluster state $\rho_{C}$ (7) can deviate from the ideal one, either because of experimental imperfections or decoherence. As the ideal state is the ground state of the nullifier, any alteration of the state results in a larger mean value of the nullifier $\textrm{Tr}[O_{k}\rho_{C}]$. To be still able to say whether a non-Gaussian cluster state was successfully prepared beyond the Gaussian ones, we will compare the nullifier mean value to the extremal value of the nullifier in Gaussian states. When the nullifier in $\rho_{C}$ overcomes this value, we can no longer say whether the state has any non-Gaussianity, at least within the framework of the non-Gaussian nullifiers. However, if the value is smaller than the Gaussian minimum, it verifies non-Gaussianity in the cluster states. This is summarized in an inequality
$\textrm{Tr}[\rho_{C}O_{k}]<\min_{G}\textrm{Tr}[\rho_{G}O_{k}],$ (10)
where $\rho_{G}$ is a general Gaussian state, and the minimization is done over the set of Gaussian states and their mixtures.
Figure 1: a Schematic illustration of a cluster state generation via $Cz$ gate and evaluation of the nullifiers $N_{1}$ and $N_{2}$ (6) from homodyne measurement results $m_{x2},m_{p1}$ and $m_{x1},m_{p2}$, respectively. b Optical implementation using beam splitter and phase shifters. c Graph of the two-node cluster state.
The natural detection of optical cluster states is homodyne measurement. Its results obey statistics represented by generalized quadratures
$X(\theta)=\cos(\theta)x+\sin(\theta)p,$ (11)
where $\theta$ is the phase given by a local oscillator. Homodyne detection under two angles $\theta=0$ and $\theta=\frac{\pi}{2}$ enables evaluation of the noise in Gaussian nullifiers. For example the measurement results $m_{p_{1}}$ and $m_{x_{2}}$ obtained by measurement of $X_{1}(\frac{\pi}{2})$ and $X_{2}(0)$ can emulate measurement of the nullifier $p_{1}-x_{2}$ as $m_{p_{1}}-m_{x_{2}}$. The approach can be generalized for non-Gaussian nullifiers if the operators $O_{n}$ can be rewritten as a finite polynomial of the form
$\expectationvalue{O_{n}}=\sum_{k}c_{k}\expectationvalue{X(\theta_{k})^{n_{k}}},$ (12)
which is a generalization of a technique previously used for the cubic nonlinear squeezing in optics Kala et al. (2022, 2025) and cubic nonlinearity based notion of nonclassicality for a mechanical oscillator Moore et al. (2019). The statistics of $X(\theta_{k})$ is approximated via the data measured with phase lock $\theta_{k}$. From them, the expectation values in (12) are estimated. The form (12) does not require measurement of cross correlations such as $\expectationvalue{X(\theta_{1})X(\theta_{2})}$.
The expression (12) offers relative freedom in the choice of the operator $O_{n}$, as every Weyl symmetric term can be written in this manner
$:x^{m}p^{n}:_{W}=\sum_{k=1}^{m+n}A_{k}X(\theta_{k})^{m+n}.$ (13)
The proof can be found in appendix A.
II Example: Photon subtracted squeezed state
Let us apply the presented non-Gaussian nullifier to a specific example of cluster state prepared from photon-subtracted squeezed states. The photon subtraction is a basic quantum operation, represented by an annihilation operator $a$ Wakui et al. (2007); Neset et al. (2024). This non-Hermitian operation is usually experimentally approximated in the following way. A small portion of the signal is reflected on a beam splitter with nearly unit transitivity and measured by a single photon detector. The realization is probabilistic, and the outcome is heralded by the successful separation of a single photon from the signal. The ideal photon-subtracted state can be rewritten as a squeezed single photon
|$\ket{\psi}_{n}=\frac{1}{\mathcal{N}}aS(r)\ket{0}=S(r)\ket{1},$ (14)
where $S(r)$ is the Gaussian squeezing operator $S(r)=\exp(\frac{r}{2}(a^{2}-a^{\dagger 2}))$. Our aim is to define a non-Gaussian nullifier that characterizes a cluster state created from photon-subtracted Gaussian squeezed states. The first step is a choice of the operator $O_{n}$ (8). The squeezed single photon state is found to be the ground state of the operator
$O_{n}=S(r)(n-1)^{2}S(-r).$ (15)
For $r=0$, $O_{n}$ consists of a number operator with shifted spectrum, such that the eigenvalue of the single photon Fock state |$\ket{1}$ equals zero. The square root then lifts all other eigenvalues to be positive.
[…]
$\begin{split}&O=X(0)^{4}(g^{4}-\frac{1}{3})+X(\frac{\pi}{2})^{4}(\frac{1}{g^{4}}-\frac{1}{3})+\frac{2}{3}(X(\frac{\pi}{4})^{4}+\\ &X(\frac{-\pi}{4})^{4})-6(X(0)^{2}g^{2}+X(\frac{\pi}{2})^{2}\frac{1}{g^{2}})+8,\end{split}$ (16)
where $g$ denotes the transformation by squeezing $S(r)$ and $g=e^{-r}$. Thus, the nullifier is well suited for the characterization of cluster states measured by homodyne measurement. The choice of $\theta_{k}$ is not unique, similarly can be used angles $0^{\circ},30^{\circ},90^{\circ}$ and $150^{\circ}$
$\begin{split}&O=S(g)[X(0)^{4}+\frac{1}{3}(\frac{16}{6}(X(30)^{4}+X(150)^{4})-3X(0)^{4}-\\ &3-\frac{1}{3}X(90)^{4})+X(90)^{4}-6X(0)^{2}-6X(90)^{2}+9]S^{\dagger}(g).\end{split}$ (17)
Next we will find the extremal value of the nullifier in Gaussian states. Due to the linearity of trace, the minimum over Gaussian states will be attained for a pure state as
$\min_{G}\textrm{Tr}[O\rho_{G}]\sim 0.611.$ (18)
The minimum was found via numerical optimization with a state
|$\ket{\psi_{G}}=S(r^{\prime})D(\alpha)\ket{0}.$ (19)
Once, the expectation value of the nullifier (15) in some state $\rho_{C}$ is below (21), we certified non-Gaussianity in the cluster state.
Another possible measurement to infer the nullifier mean value is the heterodyne measurement. The measurement can be performed by splitting the signal on a beam splitter and measuring the two arms by $x$ and $p$ homodyne. Inevitably, vacuum fluctuations enter into play from the void port of the beam splitter. The outcoming probability distribution is called the Q-function and can be obtained by convoluting the Wigner function with one unit of vacuum noise. The Q-function enables the evaluation of mean values as
[…]
where the function $O(\alpha,\alpha^{*})$ of complex variable $\alpha$ and its complex conjugate was obtained via replacing annihilation and creation operators $a\rightarrow\alpha,a^{\dagger}\rightarrow\bar{\alpha}$ in antinormally ordered expression of $O_{k}$. Due to the additional vacuum noise unit, the Gaussian minimal value will change, when evaluated from heterodyne detection to
$\min_{G}\expectationvalue{:O_{k}:_{A}}\sim 0.75.$ (21)
To derive the Gaussian extremal value, it is possible to omit the squeezing operation in (15) and include it into the optimized Gaussian state, the corresponding function of the complex amplitude $\alpha$ then yields
$O(\alpha,\bar{\alpha})=|\alpha|^{4}-5|\alpha|^{2}+4.$ (22)
Let us now analyze the behavior of the nullifier value in a more realistic scenario by considering an imperfect cluster state. The model is as follows: on a set of Gaussian squeezed states is individually performed a photon subtraction, then a Gaussian passive unitary $U$ is performed. We will consider two sources of imperfection. At first, experimentally prepared squeezed states are usually not Heisenberg limited, meaning that the amount of antisqueezing is larger compared to the squeezing. Consequently, the states are not pure. Second, the successful splitting of a single photon from the signal during photon subtraction is heralded by measurement. In practice, the heralding detector possesses only limited quantum efficiency, which can be modeled by a sequence of lossy channel with transitivity $\eta$ and an ideal detector Marek et al. (2018). Details can be found in Appendix B.
Figure 2: The value of the non-Gaussian nullifier as a function of the quantum efficiency $\eta$ of the detector heralding the photon subtraction and as a function of antisqueezing $V_{A}$[dB] of the initial state, with squeezing fixed to -2 dB.
The value of the non-Gaussian nullifier in the imperfect cluster state with varied initial antisqueezing and detector efficiency is shown in Fig. 2. Decrease in efficiency of the detector, here modeled by lossy channel with transitivity $\eta$ in front of the detector, impose higher requirements on purity of the initial squeezed state. The value of the antisqueezing has the leading effect, eventually causing the nullifier to increase over the Gaussian minimum. When the initial squeezing of -2 dB is accompanied by antisqueezing of 3.75 dB or higher (depending on efficiency of the detector), the non-Gaussianity of the cluster state is no longer verifiable within the framework of the non-Gaussian nullifier.
Apart from the imperfections of the experimental setup, the quantum properties are as well deteriorated by its inaccurate knowledge. As is schematically illustrated in Fig. 1, the knowledge of the unitary transformation producing the considered cluster state is utilized in the data processing. Let us assume, that the cluster state was prepared with a unitary operation $U$ from a set of non-Gaussian states. Additionally, let us assume that an equivalent of a different unitary transformation $U^{\prime}$ was effectively applied to evaluate the non-Gaussian nullifier. Consequently, $U^{\prime\dagger}U$ does not yield an identity operator, and $U^{\prime}\textrm{r}^{\prime}U^{\prime\dagger}$ will not address the initial statistics of the non-Gaussian states. However, the operation $U^{\prime\dagger}U$ yet not equal to identity will be again a linear transformation equivalent to an interferometer. As such, it cannot imprint any non-Gaussianity and does not embrace a risk of a false positive non-Gaussianity certification. As an illustration we include an example of the nullifier evaluation for two-mode cluster state prepared from photon-subtracted squeezed vacua that interact on a balanced beam splitter, while in the data processing a beam splitter with different transitivity $t^{\prime}=\frac{1}{\sqrt{2}}+\Delta$ was considered. The results can be seen in Fig. 3 and show the minimum of the non-Gaussian nullifier when evaluated in the true value of the transitivity and deterioration for deviation from this value by $\Delta$.
Figure 3: Value of the non-Gaussian nullifier for photon-subtracted squeezed state cluster state, which was evaluated with imperfect knowledge of the cluster preparation. The state is prepared by a balanced beam splitter, whereas the nullifier is computed for a beam splitter with transitivity $t^{\prime}=\frac{1}{\sqrt{2}}+\Delta$. The minimum can be seen when $\Delta=0$ and thus the linear operation in evaluation corresponds to an inverse of a balanced beam splitter. Three cases are shown, an ideal cluster state and then a cluster state created from photon-subtracted squeezed vacua that were subject to $10\%$ and $20\%$ losses.
Figure 4: Mean value and standard deviation of the non-Gaussian nullifier estimated from a limited dataset simulated from statistics of an ideal photon-subtracted state. The red line at $0.611$ depicts the Gaussian minimum of the nullifier. Red points correspond to initial squeezing given by $r=0.1$, blue points correspond to $r=0.2$.
The value of the non-Gaussian nullifier of an experimentally prepared cluster state can be estimated from homodyne measurement. When a phase locking is available or a data corresponding to a fixed phase can be post-selected, statistics of four generalized quadratures is sufficient in the discussed example of (16). The photon subtraction in the generation of the cluster state is probabilistic and thus requires stability of the experiment in time. Longer measurement time leads to datasets affected by experimental imperfection. To give an example, as the homodyne detection is phase sensitive, any drifting of the phase affects its precision. Therefore, we analyze the how the size of the dataset affects the precision of the nullifier estimation. We simulate random measurement values, which are drawn from an ideal theoretical probability of measuring $X(\theta)$ in an ideal cluster state. Data are further binned into $1000$ equi-width bins. In each bin, the mean value of simulated data points inside is computed together with their count. The relative frequency of the mean value appearing given by the ratio of counts in a given bin and overall number of data points then approximates the probability. Results from numerical simulation are shown in Fig. 4 for $100-10000$ measurements. The simulation was performed hundred times to determine the spread of results caused by limited sampling from the probabilities. Ten thousand measurement points for each angle are sufficient to certify the non-Gaussianity in the cluster state.
Conclusion
We defined a concept of non-Gaussian nullifiers that enables to certify a successful preparation of non-Gaussian cluster states. We showed that nullifiers defined via Weyl symmetric terms can be easily evaluated from homodyne measurement, an advantageous property as it is the natural measurement of optical cluster states. We present an example of a stabilizer for cluster state created from photon subtracted squeezed states and discuss the effect of imperfect photon subtraction and additional noise in the antisqueezed quadrature. We conclude with a note on requirements on the nullifier precise estimation.
Acknowledgement
VK and PM acknowledge grant 25-17472S of Czech Science Foundation. VK, PM and RF also acknolwedge Horizon Europe Research and Innovation Actions under Grant Agreement no. 101080173 (CLUSTEC). P.M. and R.F. acknowledge project CZ.02.01.010022_0080004649 (QUEENTEC) of EU and the Czech Ministry of Education, Youth and Sport. V.K. and R.F. acknowledges also the Quantera project CLUSSTAR (8C24003) of the Czech Ministry of Education, Youth and Sport and EU. VK acknowledges project IGA- PrF-2025-010 of the Palacký University.
Appendix A: Proof of eq. (13)
In order to prove the statement
$:x^{m}p^{n}:_{W}=\sum_{k=1}^{m+n}A_{k}X(\theta_{k})^{m+n},$ (23)
we observe that an $n$-th power of generalized quadrature equals the following sum
$X(\theta)^{n}=\sum_{k=0}^{n}[\cos(\theta)]^{n-k}[\sin(\theta)]^{k}{n\choose k}:x^{n-k}p^{k}:_{W}.$ (24)
The sum has $n+1$ Weyl-symmetric terms. In order to separate one of them in the sense of (23), we take $n$ different angles $\theta_{k}$. The coefficients $A_{k}$ can be found as a solution to a system of linear equations given by matrix
$C=\begin{pmatrix}\cos^{n}(\theta_{1})&\cos^{n}(\theta_{2})&\cdots\\ \cos^{n-1}(\theta_{1})\sin(\theta_{1})&\cos^{n-1}(\theta_{2})\sin(\theta_{2})&\cdots\\ \cos^{n-2}(\theta_{1})\sin^{2}(\theta_{1})&\cos^{n-2}(\theta_{2})\sin^{2}(\theta_{2})&\cdots\\ \vdots&\vdots&\ddots\\ \end{pmatrix},$ (25)
as
$C\textbf{A}=\textbf{d},$ (26)
where A is the vector of coefficients $A_{k}$ and d is a vector with $1$ on the position that determines which term we want to separate and otherwise equals zero. The system of linear equations has a solution, when the rank of $C$ equals the rank of the matrix with added right hand side $(C^{T}|\textbf{d})$. To this end it is sufficient to prove, that the matrix $C$ has a full rank. We start by multiplying the matrix $C$ by a diagonal matrix with diagonal
$E=\textrm{diag}(\frac{1}{\cos^{n}(\theta_{1})},\frac{1}{\cos^{n}(\theta_{2})},\frac{1}{\cos^{n}(\theta_{3})}...)$ (27)
considering different $\theta_{k}$, this will change the value of the matrix $C$, however not its rank. The resulting matrix yields
$CE=\begin{pmatrix}1&1&\cdots\\ \cot(\theta_{1})&\cot(\theta_{2})&\cdots\\ \cot^{2}(\theta_{1})&\cot^{2}(\theta_{2})&\cdots\\ \vdots&\vdots&\ddots\\ \end{pmatrix},$ (28)
and has the same rank as the matrix $C$. The rows are linearly independent and thus the matrix has a full rank.
Appendix B: Theoretical model of the imperfect photon subtracted squeezed state
We introduce two sources of imperfections in the model of the photon subtracted squeezed vacuum. At first, we replace the initial squeezed vacuum with a mixed state. This is the usual situation in experiment, where for example losses can lead to a higher antisqueezing compared to the squeezing itself. Instead of defining the mixed state, it is possible to model stat with different squeezing and antisqueezing via losses applied to a pure squeezed vacuum. The covariance matrix of the squeezed vacuum yields
$S_{0}=\begin{pmatrix}\frac{g}{2}&0\\ 0&\frac{1}{2g}\\ \end{pmatrix},$ (29)
after losses, that are modelled by a beam splitter with transitivity $t$ and reflectivity $r$, such that $t^{2}+r^{2}=1$, with the other mode being in vacuum state, the resulting density matrix is
$S_{0}=\begin{pmatrix}t\frac{g}{2}+r\frac{1}{2}&0\\ 0&t\frac{1}{2g}+r\frac{1}{2}\\ \end{pmatrix}.$ (30)
Thus the losses enable parameterization of arbitrary squeezing and antisqueezing. The squeezing or antisqueezing in dB is defined as
$\nu=10\log(\frac{V}{\frac{1}{2}}),$ (31)
where $v$ is the corresponding element of the covariance matrix. In our computation we model the state in Fock basis, defining the squeezing operation as
$S=\exp(\frac{r}{2}(a^{2}-a^{\dagger 2})),$ (32)
with
$r=\frac{\ln(g)}{2}$ (33)
and the beam splitter operation as
$U_{BS}=\exp(-r(a^{\dagger}b-ab^{\dagger})),$ (34)
where
$r=\arccos(t).$ (35)
Additionally, photon subtraction is usually provided by heralded splitting of a photon from the signal on a highly transmissive beam splitter. The heralding is done via an on/off detector, that only detects whether there is light present. This can be considered an approximation of the single photon measurement, when the intensity of the incoming light is low. The corresponding POVM element yields
$\Pi=1-\outerproduct{0}{0}.$ | . (36)
Finite quantum efficiency of the detector $\eta$ can be simulated by additional losses before an ideal detection Marek et al. (2018). Together with the losses, the corresponding POVM yields [Provaznik thesis]
$\Pi^{\prime}=1-\sum_{k=0}(1-\eta)^{k}\outerproduct{k}{k}.$ | . (37)